What is the current economic value of an inheritance that will pay $3300 to the beneficiary at the beginning of every three months for 15 years starting when the beneficiary reaches 21 years of age, 5 1/4 years from now? Assume that money can earn 6.8% compounded monthly. (Do not round intermediate calculations and round your final answer to 2 decimal places.)

  Current economic value $
Answer #1

Current Economic Value is   $ 87,787.43

Step-1:Present Value of 15 years annuity 5 1/4 years from now
Present Value =-pv(rate,nper,pmt,fv,1)
=$1,25,325.03
Where,
rate = 1.7097%
nper = 15*4 = 60
pmt = 3,300
fv = 0
Working:
Calculation of equivalent quarterly interest rate:
(1+i)^n = (1+i)^n Where,
(1+0.005667)^3 = (1+i)^1 Monthly interest rate = 6.8%/12
1.017096515 = 1+i = 0.005667
0.017096515 = i
So,
Equivalent quarterly interest rate = 1.7097%
Step-2:Calculation of current economic value
Current Economic value =-pv(rate,nper,pmt,fv)
=$87,787.43
Where,
rate = 1.7097%
nper = 5 1/ 4 years or 21 quarters = 21
pmt = 0
fv = 1,25,325.03